OpenStudy (anonymous):

Find an nth degree polynomial function with real coefficients satisfying the conditions: N=3;1and -2+3i are zeros;leading coefficients is 1

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

@FibonacciChick666

OpenStudy (anonymous):

there are a couple methods for finding a polynomial of degree 2 with zero at \(-2+3i\) a hard way, an easy way and a real real easy way take your pick

OpenStudy (fibonaccichick666):

go with sat here

OpenStudy (anonymous):

the real real easy way requires just memorizing something if the zeros are \(a+bi\) and \(a-bi\) then the quadratic is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

the fairly easy way is to take \[x=-2+3i\] and work bacwards add \(2\) \[x+2=3i\] square (carefully) to get rid of the \(i\) \[(x+2)^2=(3i)^2\\ x^2+4x+4=-9\]

OpenStudy (anonymous):

that makes the quadratic \[x^2+4x+13\] by adding 9 would have gotten the same answer with \[x^2-2\times (-2)x+(2^2+3^2)\]

OpenStudy (anonymous):

Method

OpenStudy (loser66):

we have n =3 , right?

OpenStudy (anonymous):

Yes

OpenStudy (loser66):

that is the polynomial has degree 3, right?

OpenStudy (anonymous):

Yes

OpenStudy (loser66):

so that it is in the form ax^3 + bx^2 + cx + d =0 get it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

all that writing didn't help? dang

OpenStudy (loser66):

leading coefficient =1 means a =1, oh, I am sorry @satellite73

OpenStudy (anonymous):

satellite could u help me again and accepted me as your friends

OpenStudy (loser66):

@kristinak Let read satelite's guide

OpenStudy (anonymous):

@satellite sorry I need a step by step interaction process

OpenStudy (anonymous):

Okay @Loser66

OpenStudy (anonymous):

What will be the conjugate of -2+3i

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

the conjugate of \(-2+3i\) is \(-2-3i\)

OpenStudy (anonymous):

not sure what is going on here with this question you want a polynomial with zero at \(-2+3i\) and \(1\) the quadratic that has \(-2+3i\) as a zero is \[x^2+4x+13\] so your final job is to multiply \[(x-1)(x^2+4x+13)\] to get your polynomial of degree 3

OpenStudy (anonymous):

When I set it upside it'll be (x-1)(x-2-3i)

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

you want a polynomial with real coefficients, of degree 3, not complex coefficients of degree 2

OpenStudy (anonymous):

one factor will be \((x-1)\) for sure but you also want to find the quadratic polynomial (degree 2) that has zero at \(-2+3i\) and \((-2-3i\)

OpenStudy (anonymous):

No this is just to set it up

OpenStudy (anonymous):

if you want ,(i wouldn't recommend it) you can write it as \[(x-1)(x-(-3+2i))(x-(-3-2i))\]

OpenStudy (anonymous):

if your goal is to really find the answer, a polynomial with real coefficients of degree 3, then your actual job is to multiply \[(x-1)(x^2+4x+13)\]

OpenStudy (anonymous):

That's how my teacher taught me

OpenStudy (anonymous):

Buts what's a easier way to get to the answer

OpenStudy (anonymous):

yeah i see the question i tried to explain how to get the answer before

OpenStudy (anonymous):

would you like me to try again?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok here we go you want a polynomial of degree 2 with a zero at \(-2+3i\) right?

OpenStudy (anonymous):

No degree 3

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

one of the zeros is \(1\) so one of the factors of your polynomial is \((x-1)\) is that ok?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

damn typo there one of the zeros is \(-2+3i\) the other is its conjugate \(-2+3i\) is that clear?

OpenStudy (anonymous):

ack , conjugate \(-2-3i\) that is what i meant

OpenStudy (anonymous):

so your next job is to find a polynomial of degree 2 with zeros at \(-2+3i\) and\(-2-3i\)

OpenStudy (anonymous):

Okay I got that

OpenStudy (anonymous):

lets to that

OpenStudy (anonymous):

start by setting \[x=-2+3i\] and work backwards add \(2\) to both sides, get \[x+2=3i\]

OpenStudy (anonymous):

then to get rid of the \(i\) square both sides (carefully) and get \[(x+2)^2=(3i)^2\\ x^2+4x+4=-9\]

OpenStudy (anonymous):

add \(9\) and get \[x^2+4x+13=0\] and that is your quadratic \(x^2+4x+13\)

OpenStudy (anonymous):

then to get your polynomial of degree 3, multiply \[(x-1)(x^2+4x+13)\]

OpenStudy (anonymous):

you should get \[x^3+3 x^2+9 x-13\] which i hope is one of your answers

OpenStudy (anonymous):

Let me try to do it on paper so I can show you what I get

OpenStudy (anonymous):

@satellite73 I get x^3+3x^2+9x-13

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

help me please